The data below show real wage levels at various percentiles of the wage distribution for
material-handling labourers in New Zealand.

Percentile 1975 1995
10th 4.26 2.97
50th 6.02 3.89
90th 10.44 9.07

A. Calculate w90/w50, w50/w10, and w90/w10 percentile wage differentials for both
years.
B. Describe the changes in wage inequality in this local market

C. Use the relative demand and supply model to explain the effect of increased immigrants into this labour market on relative wage and employment outcomes
Answer #1

as per HomeworkLib rules answered first 2 parts

Solution:

Given information of real wage levels at various percentiles of the wage distribution for material-handling labourers in New Zealand.

A. Calculation of  w90/w50, w50/w10, and w90/w10 percentile wage differentials for both
years are:

Year 1975:

w90/w50= 10.44/6.02= 1.734

w50/w10= 6.02/4.26= 1.413

w90/w10= 10.44/ 4.26= 2.450

for year 1995:

w90/w50= 9.07/3.89 = 2.331

w50/w10= 3.89/2.97 = 1.309

w90/w10= 9.07/2.97 = 3.053

B. The changes in wage inequality in this local market is:

from above calculations we can see that ,

Wage inequality has grown from 1975 to 1995 for w90/w50 and w90/w10 and has shrinked for w50/w10.